OpenStudy (anonymous):

write the quadratic function f(x) = 3x^2 -12x +1 in the form f(x) = 3(x-h)^2+k

OpenStudy (anonymous):

Complete the square.

OpenStudy (anonymous):

\[ 3x^2 -12x +1=3(x^2 -4x) +1 \]

OpenStudy (anonymous):

You complete the square on \(x^2-4x\).

OpenStudy (anonymous):

or cheat, which is what i prefer

OpenStudy (anonymous):

\[ x^2+bx = \left(x+\frac b 2\right)^2-\left(\frac b 2\right)^2 \]

OpenStudy (anonymous):

cheating method saves time on test or quiz vertex of \(ax^2+bx+c\) is always \(-\frac{b}{2a}\) in your case it is \(-\frac{-12}{2\times 3}=2\) so you know it will look like \[3(x-2)^2+k\] and you can find \(k\) by replacing \(x\) by \(2\) in the original expression